### v3

# Refa: Reference Free Alignment for Multi-preference Optimization

Abstract
We introduce Refa, a family of reference-free alignment methods that optimize over multiple user preferences while enforcing fine-grained length control. Our approach integrates deviation-based weighting to emphasize high-quality responses, length normalization to prevent trivial short-response solutions, and an EOS-probability regularizer to mitigate dataset-induced brevity biases. Theoretically, we show that under the Uncertainty Reduction with Sequence Length Assertion (URSLA) framework, naïve length normalization can still incentivize length-based shortcuts. In contrast, Refa corrects these subtle incentives, guiding models toward genuinely more informative and higher-quality outputs. Empirically, Refa achieves a new state-of-the-art among reference-free alignment methods, generating richer responses that align more closely with human preferences. Notably, Refa improves performance on the AlpacaEval2 benchmark, achieving a 26.6% Length-Controlled Win Rate (LC-WR) and 24.2% Win Rate (WR).
Machine Learning, ICML
1 Introduction
As large language models (LLMs) find widespread deployment in real-world applications—such as customer support, educational tutoring, and medical advice—ensuring that their outputs adhere to human values, quality standards, and usage guidelines has become a fundamental research objective (Ouyang et al., 2022; Liu et al., 2023; Rafailov et al., 2024). Alignment methods, notably Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020) and Direct Preference Optimization (DPO) (Rafailov et al., 2024; Xu et al., 2024b), have proven essential for guiding models toward producing safer, more helpful content that meets user expectations.
While many alignment techniques rely on a reference model (Ouyang et al., 2022; Christiano et al., 2017), there is growing interest in reference-free alignment approaches (Cui et al., 2023; Liu et al., 2024; Zhou et al., 2023; Wang et al., 2024). By directly optimizing from scalar rewards or scores, these methods sidestep complex ratio modeling and exploit richer information from scenarios where multiple positive and negative responses per query are available. Utilizing such diverse signals has been shown to improve downstream performance (Cui et al., 2023; Yang et al., 2024), indicating that reference-free alignment can offer more robust and transferable improvements in language model capabilities.
Traditional alignment formulations often adopt a pairwise stance, contrasting a single preferred response against a single dispreferred one (Rafailov et al., 2024; Xu et al., 2024b). However, real-world user preferences are more nuanced. Users may have multiple acceptable responses to a given query, as well as several suboptimal or even unacceptable ones. This observation motivates multi-preference optimization, an approach that jointly considers entire sets of positive and negative responses (Gupta et al., 2024; Chen et al., 2024), enabling models to align with a richer, more representative distribution of desirable outputs.
Yet, multi-preference optimization introduces new challenges, particularly the model’s tendency to exploit dataset-induced biases toward brevity. Humans often produce concise training examples due to cognitive and economic constraints, inadvertently suggesting shorter answers as the norm. In contrast, users at inference time frequently prefer richer, more detailed explanations—especially for complex tasks requiring extended reasoning. Without careful design, a model may prematurely terminate responses, undermining genuine quality improvements (Holtzman et al., 2019; Wolf et al., 2023). Controlling length via length-regularization is thus critical. Merely normalizing by length, à la Simpo(Meng et al., 2024) or RRHF(Yuan et al., 2023), is insufficient as length biases may still persist due to differences in response lengths. In this form of reward hacking, the model may learn to rely on length as a spurious correlate to reward instead of answer semantics. Instead of relying on brevity as a shortcut, we should realign training incentives to encourage more detailed outputs, ensuring that improvements stem from genuinely higher-quality responses rather than truncated answers that exploit training biases.
In this work, we propose Refa, a family of reference-free multi-preference alignment strategies that tackle these issues. Our approach integrates deviation-based weighting, ensuring that responses far above or below average quality receive proportionally more influence (Gupta et al., 2024), and introduces length normalization coupled with an EOS-probability regularizer. These techniques mitigate dataset-induced biases and encourage models to produce the richer, detailed answers often required in real-world scenarios. Moreover, we provide theoretical insights and gradient-level analyses of our loss function with respect to both response probabilities and length, offering guarantees on the fixed points of the InfoNCA loss and our Refa objective. By combining strong theoretical grounding with empirical validation on benchmark datasets, we demonstrate that our Refa-dynamic formulation achieves state-of-the-art results, advancing the alignment techniques toward more robust, controllable, and preference-aware LLM training.
1.1 Our Contributions
In this work, we develop Refa, a family of reference-free, multi-preference alignment methods that integrate length normalization, deviation-based weighting, and an EOS-probability regularizer to encourage richer responses. Our key contributions include:
  1. 1.
    Algorithmic Novelty: We introduce a new Refa-dynamic variant that leverages deviation-based weighting and a carefully designed length-regularizer. This combination resolves the subtle incentives that previous methods faced, ensuring that improvements in alignment arise from genuinely higher-quality responses rather than trivial length manipulations.
  2. 2.
    Theoretical Insights: We formalize the Uncertainty Reduction with Sequence Length Assertion (URSLA), a conjecture stating that, for certain subsets of responses, increasing sequence length reduces per-token uncertainty. Under URSLA, we prove that naïve length normalization can inadvertently incentivize shorter negative responses, illuminating why additional constraints (like EOS probability control) are essential. Furthermore, we offer theoretical insights relating to the stationary point, and gradient updates of the various preference-optimization algorithms.
  3. 3.
    State-of-the-art Results: Our experiments on the AlpacaEval v2 benchmark demonstrate that Refa significantly outperforms strong baselines. Compared to the strongest multi-preference baseline (InfoNCA) and the best reference-free baseline (SimPO), our Refa-dynamic (p=2) formulation achieves the highest length-controlled and raw win rates (24.6% LC-WR and 23.0% WR for Mistral-Base (7B)), validating the effectiveness of our approach in producing richer, higher-quality responses.
Related Works:
We provide a detailed description of our related work in Appendix A covering other preference optimization, reference-free alignment as well as multi-preference optimization approaches.
2 Notation and Preliminaries
In this section, we establish the notation and fundamental concepts used throughout the paper. We consider a reference-free multi-preference alignment problem where each query can be paired with multiple responses of varying quality. The objective is to shape the policy distribution so that it assigns higher probability mass to higher-quality responses, while pushing the probabilities of lower-quality ones closer to zero.
Queries and Responses
Let $\mathcal{X}$ denote the space of possible queries. For each query $x\in\mathcal{X}$, consider a set of candidate responses $\{y_{i}\}_{i=1}^{K}\subseteq\mathcal{Y}$. Each response $y_{i}$ is a sequence of tokens whose length is denoted by $|y_{i}|$. This length $|y_{i}|$ plays a crucial role in later sections, where we introduce length normalization to prevent trivial short-response biases.
Policy Model
We consider a policy model $\pi_{\theta}(y\mid x)$, parameterized by $\theta$, that assigns probabilities to candidate responses for each query. In particular, $\pi_{\theta}(y\mid x)=\exp(\log\pi_{\theta}(y\mid x))$, where $\log\pi_{\theta}(y\mid x)$ is the log-probability of $y$. Throughout this work, we focus on a reference-free scenario, meaning we do not rely on a fixed baseline distribution $\mu(y\mid x)$.
Rewards and Mean Rating
Each response $y_{i}$ to a query $x$ is associated with a scalar reward $r_{i}=r(x,y_{i})$, which can be interpreted as a measure of the response’s quality. Let $\bar{r}=\tfrac{1}{K}\sum_{i=1}^{K}r_{i}$ denote the mean reward across the candidate responses. We define the positive set $Y^{+}=\{y_{i}\mid r_{i}>\bar{r}\}$ consisting of responses with above-mean ratings, and the negative set $Y^{-}=\{y_{i}\mid r_{i}\leq\bar{r}\}$ composed of responses at or below the mean rating.
Target Distribution and Multi-Preference Alignment
Some alignment methods within the Refa family, such as the InfoNCA variant, incorporate a target distribution derived from the rewards. This involves introducing a temperature parameter $\alpha>0$ and defining
$p_{i}^{\text{target}}=\frac{e^{\alpha r_{i}}}{\sum_{j=1}^{K}e^{\alpha r_{j}}}.$
The Refa-InfoNCA method encourages the model distribution to align with $p_{i}^{\text{target}}$, thereby treating all responses according to their relative quality as encoded by the rewards. In contrast, other proposed variants of Refa directly seek to drive down the probability of negative responses without relying on this target distribution.
Deviations and Weighting
To capture nuanced differences among responses beyond simple binary classification as positive or negative, we consider deviation-based weighting. Let $\Delta S_{i}=r_{i}-\bar{r}$ denote the deviation of a response’s reward from the mean rating. Assigning weights based on $\Delta S_{i}$, for example $w_{i}=e^{\alpha\Delta S_{i}}$, allows us to emphasize highly positive responses more strongly and penalize significantly negative ones more intensely.
Length Normalization and EOS Regularization
Language models may find certain token sequences easier to predict, causing shorter responses to be disproportionately favored due to their higher probability. To prevent trivial solutions such as always producing very short responses, we apply length normalization. Instead of using $\log\pi_{\theta}(y\mid x)$ directly, we employ a length-normalized quantity
$\overline{\log\pi_{\theta}}(y\mid x)=\frac{1}{|y|}\sum_{\text{tokens }t\in y}\log P_{\theta}(t\mid\text{context}),$
ensuring that the measure of quality is not conflated with response brevity. Furthermore, to maintain control over response lengths, we introduce an end-of-sequence (EOS) regularizer that penalizes placing excessive probability mass on early termination tokens. This regularizer provides finer control over response length and ensures that improved alignment does not emerge solely from truncating responses.
Hyperparameters
Several hyperparameters regulate the behavior of our reference-free multi-preference alignment framework. The parameter $\alpha$ influences how rating differences translate into distributional adjustments or weighting schemes. The parameter $\beta$ acts as an inverse temperature that sharpens or smooths distributions derived from the model’s log-probabilities. The parameter $\gamma$ controls how strongly negative responses are penalized relative to positive ones. Finally, the parameter $\lambda$ weights the length or EOS-based regularizer, thereby adjusting the extent to which sequence length considerations factor into the optimization. By tuning $(\alpha,\beta,\gamma,\lambda)$, it becomes possible to adapt the training objective to various tasks, data distributions, and desired response characteristics, ensuring that improvements in alignment reflect enhancements in response quality.
3 Methodology: Deriving the Refa Loss Function
In this section, we present a step-by-step derivation of our Refa loss function, starting from the InfoNCA formulation, removing reference models, introducing multi-preference sets, incorporating deviation-based weighting, length normalization, and finally addressing length-related biases through an EOS-regularization mechanism. For detailed proofs, intermediate steps, and nuanced discussions, we refer the reader to the appendices: Appendix B through Appendix G.4.
3.1 Preliminary Setup and Notation
We consider a query $x\in\mathcal{X}$ and a set of candidate responses $\{y_{i}\}_{i=1}^{K}\subseteq\mathcal{Y}$. Each response $y$ is a sequence of tokens, and we define the full context (e.g., prompt or conversation history) under which $y$ is generated as $\text{context}:=(x)$. Our policy model $\pi_{\theta}(y\mid\text{context})$ assigns probabilities to responses conditioned on this context. Throughout this section, the term “context” encompasses all relevant conditioning information available to the model, including the query $x$.
We assign to each response $y$ a scalar reward $r_{y}$ and define the mean reward $\bar{r}:=\tfrac{1}{K}\sum_{i=1}^{K}r_{y_{i}}$. Based on this, we partition responses into:
$Y^{+}$ =$=\{y\mid r_{y}>\bar{r}\},$ (1)
$Y^{-}$ =$=\{y\mid r_{y}\leq\bar{r}\}.$ (2)
To handle variable-length sequences, we define the length-normalized log-likelihood of a response:
$\overline{\log\pi_{\theta}}(y\mid x)\;:=\;\frac{1}{|y|}\sum_{\text{token }t\in y}\log P_{\theta}(t\mid\text{context}).$ (3)
This ensures that per-token quality is measured independently of the response length.
3.2 From InfoNCA to Reference-Free Alignment
InfoNCA with a Reference Model:
The InfoNCA approach (Chen et al., 2024) originally assumes access to a fixed reference model $\mu(y\mid x)$ and aligns the ratio $\frac{\pi_{\theta}(y\mid x)}{\mu(y\mid x)}$ to a distribution induced by rewards. In its original form, given a target distribution $p_{i}^{\text{target}}$ derived via a softmax over rewards (Appendix B), the InfoNCA loss can be expressed as:
$L_{\text{InfoNCA}}(\theta)$ $=-\sum_{i=1}^{K}p_{i}^{\text{target}}\log p_{i}^{\text{model}},$ (4)
where
$p_{i}^{\text{model}}$ $=\frac{\frac{\pi_{\theta}(y_{i}\mid x)}{\mu(y_{i}\mid x)}}{\sum_{j=1}^{K}\frac{\pi_{\theta}(y_{j}\mid x)}{\mu(y_{j}\mid x)}}.$ (5)
Removing the Reference Model:
Eliminating the reference distribution $\mu(y\mid x)$ yields a simpler, reference-free setting (Appendix C). Without $\mu$, the model distribution directly becomes:
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}\mid x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}\mid x)}.$ (6)
However, this direct approach has several issues (Appendix D): it can promote low-quality responses, create contradictory optimization signals among high-quality responses, and fail to distinguish subtle differences in reward quality.
3.3 Multi-Preference Alignment via Positive and Negative Sets
To address these weaknesses, we consider a multi-preference formulation (Appendix E): rather than focusing on pairwise or single-response comparisons, we split responses into sets $Y^{+}$ and $Y^{-}$. We want to collectively increase probabilities for all $y\in Y^{+}$ and decrease probabilities for all $y\in Y^{-}$. A starting point is:
$L(\theta)=-\log\frac{\sum_{y\in Y^{+}}\pi_{\theta}(y\mid x)}{\sum_{y\in Y^{+}\cup Y^{-}}\pi_{\theta}(y\mid x)}.$ (7)
This ensures that to decrease $L(\theta)$, the model must raise the relative mass of $Y^{+}$ and reduce that of $Y^{-}$. Yet, this basic version treats all responses in $Y^{+}$ or $Y^{-}$ equally, ignoring finer reward gradations.
3.4 Deviation-Based Weighting and Hyperparameters
To incorporate the full spectrum of quality differences, we introduce a deviation-based weighting (Appendix D) defined by:
$\Delta S_{y}$ $=r_{y}-\bar{r},$ (8)
$w_{y}$ $=e^{\alpha\Delta S_{y}},$ (9)
where $\alpha>0$ controls how strongly reward deviations influence weighting. Such a weighting also induces a curriculum in training, helping the optimization process focus on the high weight examples first, and then the median examples.
We also introduce inverse-temperature parameters $\beta$ and a negative-set weighting factor $\gamma$. The inverse temperature $\beta$ sharpens or smooths the distribution, and $\gamma>0$ controls how heavily to penalize the negative set. Incorporating these parameters and weights, we have:
$L(\theta)$ $=-\log\frac{P^{+}}{P^{+}+\gamma P^{-}}$
where
$P^{+}$ $=\sum_{y\in Y^{+}}e^{\beta(\log\pi_{\theta}(y\mid x)+\alpha\Delta S_{y})},$
$P^{-}$ $=\sum_{y\in Y^{-}}e^{\beta(\log\pi_{\theta}(y\mid x)+\alpha\Delta S_{y})}$ (10)
This aligns the model distribution such that high-deviation (high-quality) responses in $Y^{+}$ gain more probability, and low-quality responses in $Y^{-}$ are penalized more intensely.
3.5 Length Normalization and EOS Probability Regularization
A key insight from Appendix G is that without length normalization, shorter responses gain an unfair advantage as they have fewer tokens to predict. To counter this, we replace $\log\pi_{\theta}(y\mid x)$ with its length-normalized counterpart $\overline{\log\pi_{\theta}}(y\mid x)$ as defined in Eq. equation 3:
$L(\theta)$ $=-\log\frac{\overline{P^{+}}}{\overline{P^{+}}+\gamma\overline{P^{-}}}$
where
$\overline{P^{+}}$ $=\sum_{y\in Y^{+}}e^{\beta(\overline{\log\pi_{\theta}}(y\mid x)+\alpha\Delta S_{y})},$
$\overline{P^{-}}$ $=\sum_{y\in Y^{-}}e^{\beta(\overline{\log\pi_{\theta}}(y\mid x)+\alpha\Delta S_{y})}$ (12)
Finally, due to dataset-induced biases and the risk of models exploiting length shortcuts (Appendix G.4), we introduce an EOS-regularization term $\mathcal{R}(\theta)$. This term encourages or discourages early termination, ensuring that the model does not trivially shorten responses to reduce loss. The final Refa objective is:
$L_{\textsc{Refa}}(\theta)$ $=-\log\frac{\overline{P^{+}}}{\overline{P^{+}}+\gamma\overline{P^{-}}}\;+\;\mathcal{R}(\theta).$ (13)
4 Algorithm
The following procedure outlines how to train a model with the proposed Refa loss. We assume a dataset of queries and associated responses with scalar rewards, as well as known hyperparameters.
Algorithm 1 REFA Training Procedure
1:  Input: Dataset $\{(x,\{y_{i},r_{i}\}_{i=1}^{K})\}_{n=1}^{N}$, learning rate $\eta$, hyperparameters $\alpha,\beta,\gamma,\lambda$
2:  repeat
3:     Sample a batch of queries $\{x\}$ and their responses $\{y_{i},r_{i}\}$
4:     for all query $x$ in the batch [^1] For brevity, we omit the data index superscript $n$ throughout the algorithm. do
5:        Compute mean reward: $\bar{r}\leftarrow\frac{1}{K}\sum_{i=1}^{K}r_{i}$
6:        Split responses:$Y^{+}\leftarrow\{y_{i}\mid r_{i}>\bar{r}\}$, $Y^{-}\leftarrow\{y_{i}\mid r_{i}\leq\bar{r}\}$
7:        Compute length-normalized log-probs: $\overline{\log\pi_{\theta}}(y_{i}\mid x)$
8:        Compute deviation-based weights: $w_{i}\leftarrow\exp(\alpha(r_{i}-\bar{r}))$
[…]
10:        Compute EOS regularizer $\mathcal{R}(\theta)$
11:        Compute loss:
12:        $L_{\textsc{Refa}}(\theta)\leftarrow-\log\left(\frac{\overline{P^{+}}}{\overline{P^{+}}+\gamma\overline{P^{-}}}\right)+\mathcal{R}(\theta)$
13:     end for
14:     Compute batch loss: $L_{\text{batch}}(\theta)\leftarrow\frac{1}{|\text{batch}|}\sum_{n}L_{\textsc{Refa}}(\theta)$
15:     Update parameters: $\theta\leftarrow\theta-\eta\nabla_{\theta}L_{\text{batch}}(\theta)$
16:  until convergence
Explanation:
The algorithm implements a contrastive learning approach for response alignment. At each iteration, it samples query-response pairs (line 2) and computes mean rewards (line 4) to partition responses into above-mean ($Y^{+}$) and below-mean ($Y^{-}$) sets (line 5). Length-normalized log-probabilities (line 6) prevent degenerate solutions, while deviation-based weights (line 7) sharpen the preference signal.
The core learning objective aggregates weighted probabilities through partition functions $\overline{A}$ and $\overline{B}$ (lines 8-9), with an EOS regularizer $\mathcal{R}(\theta)$ (line 10) to control response termination. The final loss (line 11) drives gradient updates to optimize the model’s policy distribution (lines 13-14).
5 Theoretical Analysis
We now highlight the central theoretical insights derived from our reference-free multi-preference alignment framework. Starting from a reference-free version of the InfoNCA objective, we identify fundamental issues—such as encouraging low-quality responses and creating contradictory optimization signals—when dealing with multiple responses of varying quality. We then show how length normalization and additional regularizers (like EOS probability control) address subtle incentives that arise due to the relationship between sequence length and token-level uncertainty.
5.1 Single-Term Objectives and Their Gradients
In a reference-free setting, the InfoNCA loss reduces to a cross-entropy objective over multiple candidate responses. However, analyzing it at the granularity of single responses (i.e., single-term objectives) reveals how the model’s gradient signals can inadvertently promote low-quality outputs.
Lemma 5.1 (Gradient of a Single-Term Objective).
(Originally stated in Appendix C.2) Let $\ell_{i}(\theta):=-p_{i}^{\text{target}}\log p_{i}^{\text{model}}$, where
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}$
and $p_{i}^{\text{target}}>0$. Suppose $\pi_{\theta}(y_{j}|x)>0$ for all $j$. Then:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}=-p_{i}^{\text{target}}\left(\frac{1}{\pi_{\theta}(y_{i}|x)}-\frac{1}{\sum_{j}\pi_{\theta}(y_{j}|x)}\right),$
and for each $j\neq i$,
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}=p_{i}^{\text{target}}\frac{1}{\sum_{k}\pi_{\theta}(y_{k}|x)}.$
Proof Sketch:
1. Express $\ell_{i}(\theta)$ in terms of $\log\pi_{\theta}(y_{i}|x)$ and $\log(\sum_{j}\pi_{\theta}(y_{j}|x))$. 2. Differentiate w.r.t. $\pi_{\theta}(y_{i}|x)$ using the chain rule: $\partial_{\pi_{\theta}(y_{i}|x)}\log\pi_{\theta}(y_{i}|x)=1/\pi_{\theta}(y_{i}|x)$ and $\partial_{\pi_{\theta}(y_{i}|x)}\log(\sum_{j}\pi_{\theta}(y_{j}|x))=1/\sum_{j}\pi_{\theta}(y_{j}|x)$. 3. For $j\neq i$, only the denominator $\sum_{j}\pi_{\theta}(y_{j}|x)$ affects the gradient. (For full proof see Appendix C.2.)
Lemma 5.2 (Directional Influence of a Single-Term Objective).
(Originally stated in Appendix C.2) Under the same assumptions as Lemma 5.1, minimizing $\ell_{i}(\theta)$ encourages increasing $\pi_{\theta}(y_{i}|x)$ (especially if initially small) and decreasing $\pi_{\theta}(y_{j}|x)$ for $j\neq i$.
Proof Sketch:
1. From Lemma 5.1, note that if $\pi_{\theta}(y_{i}|x)$ is small, then $\frac{\partial\ell_{i}}{\partial\pi_{\theta}(y_{i}|x)}<0$, so increasing $\pi_{\theta}(y_{i}|x)$ reduces $\ell_{i}(\theta)$. 2. For $j\neq i$, $\frac{\partial\ell_{i}}{\partial\pi_{\theta}(y_{j}|x)}>0$, so decreasing $\pi_{\theta}(y_{j}|x)$ reduces $\ell_{i}(\theta)$. (For full proof see Appendix C.2.)
Implication:
These lemmas show that when optimizing each response independently, even low-quality responses receive upward pressure if their target weight $p_{i}^{\text{target}}$ is nonzero. Moreover, multiple good responses compete with each other, causing contradictory optimization signals.
5.2 Length as a Shortcut: Probability Inflation and Length-Induced Certainty
Without length normalization, the model can exploit shorter sequences to artificially inflate their probabilities. Even with length normalization, subtle effects arise from the interplay between sequence length and token-level uncertainty.
Lemma 5.3 (Length-Induced Probability Inflation).
(Originally stated in Appendix F, Lemma: Length-Induced Probability Inflation) Shorter responses typically accumulate less negative log-probability because they have fewer tokens. Thus, holding per-token quality constant, shorter responses achieve higher overall probability, incentivizing the model to prefer brevity over meaningful improvement.
Proof Sketch:
1. Consider two responses $y_{1},y_{2}$ with similar per-token likelihoods. 2. If $|y_{1}|<|y_{2}|$, then $\log\pi_{\theta}(y_{1}|x)=-|y_{1}|c$ and $\log\pi_{\theta}(y_{2}|x)=-|y_{2}|c$ for some average cost $c>0$. 3. Since $|y_{1}|<|y_{2}|$, $e^{-|y_{1}|c}>e^{-|y_{2}|c}$, so $\pi_{\theta}(y_{1}|x)>\pi_{\theta}(y_{2}|x)$. (For full proof see Appendix F.)
5.3 Uncertainty Reduction with Sequence Length Assertion (URSLA)
We introduce a conjecture that articulates how model uncertainty changes with response length. Let $y$ denote a response (sequence of tokens) and $\overline{LP}(y)$ be the average per-token negative log-probability of $y$ under the model $\pi_{\theta}$. i.e. $\overline{LP}(y)=P_{\theta}(y)/\operatorname{len}(y)$. Then, lower $\overline{LP}(y)$ corresponds to higher model confidence at the per-token level.
Conjecture 5.4 (Uncertainty Reduction with Sequence Length Assertion (URSLA)).
There exists a non-empty subset of responses $\mathcal{Y}$ such that for all $y\in\mathcal{Y}$, increasing the length of $y$ reduces the average per-token negative log-probability $\overline{LP}(y)$. Formally:
$\frac{\partial\overline{LP}(y)}{\partial\operatorname{len}(y)}<0\quad\text{forall }y\in\mathcal{Y}.$
In other words, for these responses, as sequences grow longer, they become progressively easier for the model to predict, lowering the per-token uncertainty.
Limitations of Conjecture: This conjecture aligns with empirical observations: once the model establishes a coherent semantic trajectory, subsequent tokens become more predictable. However, URSLA may fail in scenarios where the model encounters domain shifts, complex reasoning, or adversarial inputs. Domain-specific empirical verification is needed to confirm when this conjecture holds. (See Appendix G for a detailed discussion.)
Figure 1: Empirical analysis for URSLA on Benchmark dataset.
5.4 Consequences of URSLA for Negative Responses
Under many preference optimization objectives, certain responses—labeled as “negative”—are penalized if they have too high a probability. To reduce their probability, one can equivalently seek to increase $\overline{LP}(y)$ for these responses. If URSLA holds, longer sequences are inherently more predictable (lower $\overline{LP}(y)$), so to raise $\overline{LP}(y)$ for a negative response, the model may choose to shorten it.
Lemma 5.5 (Length Reduction for Negative Responses).
(Originally stated in Appendix G, Lemma: Length Reduction for Negative Responses) Suppose we have a loss function $L(\theta)$ that penalizes high-probability negative responses. Increasing the average log-perplexity $\overline{LP}(y)$ for these negative responses decreases $L(\theta)$. Under the URSLA Conjecture 5.4, because longer sequences have lower $\overline{LP}(y)$, the model can increase $\overline{LP}(y)$ by shortening these responses. Formally, if:
$\frac{\partial L(\theta)}{\partial\operatorname{len}(y)}>0\text{ for all }y\in\mathcal{Y}^{-},$ (14)
then minimizing $L(\theta)$ encourages decreasing $\operatorname{len}(y)$ for negative responses $y\in\mathcal{Y}^{-}$.
Proof (Sketch): If the objective penalizes a negative response $y$, it aims to lower $\pi_{\theta}(y|x)$. Equivalently, it must increase $\overline{LP}(y)$. Under URSLA, $\overline{LP}(y)$ decreases with length, so making $y$ shorter increases $\overline{LP}(y)$, aiding in lowering $\overline{\log\pi_{\theta}}(y|x)$, which is the expression in the second term in the denominator of the loss function. Thus, reducing length is a shortcut to achieving the training objective, which implies a positive partial derivative of $L(\theta)$ w.r.t. response length for negative responses. This creates an incentive to shorten negative responses to reduce the loss.
Implication:
Even with length normalization, subtle incentives remain. If we must penalize certain negative responses, the model can exploit length manipulation to achieve the required uncertainty. This suggests that length normalization alone is insufficient for controlling response quality and justifies adding further constraints, such as EOS probability regularizers.
Summary of Theoretical Insights
Our analysis reveals fundamental limitations of reference-free InfoNCA objectives, which can promote low-quality responses, create counterproductive competition, and enable length-based shortcuts—even under length normalization. These insights motivate integrating deviation-based weighting, multi-preference sets, and explicit length control to better align model outputs with user preferences.
6 Experiments
Method Mistral-Base (7B) Llama-3-Base (8B)
AlpacaEval 2 Arena-Hard MT-Bench AlpacaEval 2 Arena-Hard MT-Bench
LC (%) WR (%) WR (%) GPT-4 LC (%) WR (%) WR (%) GPT-4
SFT [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 8.4 6.2 1.3 6.3 6.2 4.6 3.3 6.6
RRHF [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 11.6 10.2 5.8 6.7 12.1 10.1 6.3 7.0
SLiC-HF [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 10.9 8.9 7.3 7.4 12.3 13.7 6.0 7.6
DPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 15.1 12.5 10.4 7.3 18.2 15.5 15.9 7.7
IPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 10.8 9.0 7.5 7.0 14.4 14.2 17.8 7.4
CPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 9.8 8.9 6.7 6.9 10.8 8.1 5.8 7.4
KTO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 13.1 9.1 5.6 6.8 14.2 12.4 12.5 7.8
ORPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 14.7 12.2 7.0 7.0 12.2 10.6 10.8 7.6
R-DPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 17.4 12.8 10.5 7.0 17.6 14.4 17.2 7.5
SWEPO [^2] These numbers are taken directly from the paper Swepo: Simultaneous Weighted Preference Optimization for Group Contrastive Alignment (Gupta et al., 2024) 20.3 14.9 12.8 7.3 20.1 15.6 18.5 7.8
SimPO [^1] These are taken directly from the paper Simpo: Simple Preference Optimization with a Reference-Free Reward (Meng et al., 2024) 21.5 20.8 16.6 7.3 22.0 20.3 23.4 7.7
Refa-InfoNCA 20.4 19.1 14.2 7.1 22.8 17.1 20.2 7.5
Refa-InfoNCA (dynamic) 19.0 18.4 13.9 7.1 21.4 21.5 25.1 7.6
Refa-1-vs-k (p=0) 20.4 19.1 14.6 7.1 25.5 21.0 26.5 7.5
Refa-1-vs-k (p=1) 20.1 17.8 12.6 7.1 26.5 21.3 24.5 7.5
Refa-1-vs-k (p=2) 21.7 19.6 14.4 7.1 24.0 18.1 21.7 7.4
Refa-dynamic (p=0) 22.8 20.4 15.7 7.3 25.6 23.1 25.4 7.8
Refa-dynamic (p=1) 24.3 22.4 16.2 7.3 26.3 23.2 22.8 7.8
Refa-dynamic (p=2) 24.6 23.0 16.6 7.4 26.6 24.2 23.7 7.8
Table 1: Comparison of various preference optimization baselines on AlpacaEval, Arena-Hard and MT-Bench benchmarks. LC-WR represents length-controlled win rate, and WR represents raw win rate. Best results are in bold, second-best are underlined. Our method (Refa) achieves SOTA performance across all metrics, with different variants achieving either best or second-best results consistently.
Model and Training Settings:
For our experiments, we utilized the Ultrafeedback Dataset (Cui et al., 2023), an instruction-following benchmark annotated by GPT-4. This dataset consists of approximately 64,000 instructions, each paired with four responses generated by different language models. GPT-4 assigned scalar rewards on a 0-to-10 scale for each response, which prior research has shown to correlate strongly with human annotations. This establishes GPT-4 ratings as a reliable and cost-efficient alternative to manual feedback.
In our broader framework, we first trained a base model (mistralai/Mistral-7B-v0.1) on the UltraChat-200k dataset to obtain an SFT model. This SFT model, trained on open-source data, provides a transparent starting point. Subsequently, we refined the model by performing preference optimization on the UltraFeedback dataset. Once fine-tuned, the model was used for alignment. This two-step process ensures the model is well-prepared for tasks.
In our experiments, we observed that tuning hyperparameters is critical for optimizing the performance of all offline preference optimization algorithms, including DPO, SimPO, Refa-InfoNCA, Refa-1-vs-all and Refa-dynamic. Hyper-parameter selection is very crucial for these.
For Refa-InfoNCA, we found that setting the beta parameter in the range of 2.0 to 2.5 consistently yields strong performance. Similarly, for Refa-1-vs-all and Refa-dynamic, optimal results were achieved with beta values between 2.0 and 4.0, while tuning the gamma parameter within the range of 0.7 to 1.4 further improved performance. These observations highlight the importance of systematic hyperparameter tuning to achieve reliable outcomes across diverse datasets.
Evaluation Benchmarks
We evaluate on MT-Bench (80 questions across 8 categories), AlpacaEval 2 (805 questions from 5 datasets), and Arena-Hard (500 technical problems). For AlpacaEval 2, we report win rate (WR) and length-controlled win rate (LC). For Arena-Hard, we report win rate against a baseline. For MT-Bench, we use GPT-4-Preview-1106 as judge. We provide full details of our evaluation benchmarks as well as baselines in Appendix H.
6.1 Experimental Results
Figure 2: EOS Probability Analysis Across Training Steps: Demonstrating how EOS probability regularization reduces premature EOS token generation as optimization progresses for mistral and llama.
Figure 3: Impact of $\lambda$ on Mistral-Base (7B) Performance: a) Average Response Length, b) Length-Controlled Win Rate (LC-WR), and c) Win Rate (WR).
Figure 4: Effect of penality-ratio $\gamma$ on Mistral-Base (7B) on AlpacaEval2 Benchmark.
Refa-dynamic outperforms existing preference baselines:
We conducted an extensive comparison of the Refa-Dynamic framework against various preference optimization methods using the Mistral-7B model, as summarized in Table 1. The results demonstrate that Refa-Dynamic consistently outperforms all baselines across the primary evaluation metrics. Specifically, Refa-Dynamic achieves improvement of 3.1% on LC-WR and 2.2% on WR over SimPO, which is the second-best baseline. Compared to other competitive approaches such as R-DPO and InfoNCA, Refa-Dynamic delivers significant gains, showcasing its ability to handle diverse preferences. These results highlight the efficacy of REFA-Dynamic’s reference-free optimization in addressing alignment biases and enhancing response quality, reinforcing its state-of-the-art performance in preference optimization.
Effect of EOS Probability Regularization on Refa-dynamic:
[…]
Effect of Penalty-Ratio $\gamma$ on Refa-dynamic:
To investigate the impact of the penalty-ratio parameter $\gamma$ on the performance of ReFa-Dynamic, we evaluate two key metrics: LC-WR (%) and WR (%) across varying values of $\gamma$. As shown in Figure 4, LC-WR increases initially, peaking at $\gamma=3$, before slightly declining. This trend indicates that a moderate penalty ratio balances positive and negative response contributions effectively, improving alignment consistency. Conversely, WRshows a decreasing trend as $\gamma$ increases beyond 3, suggesting that overly penalizing negative responses may compromise reward alignment.
The dual-axis plot highlights the trade-off: while higher $\gamma$ values enhance LC-WR by enforcing stricter penalties on negative responses, they may inadvertently reduce WR%. Key findings underline the importance of selecting an optimal $\gamma$ to maintain a balance between alignment consistency and reward fidelity.
Refa-InfoNCA (1-vs-all) vs Refa-InfoNCA (dynamic) vs Refa-InfoNCA
We evaluated the performance of Refa-InfoNCA and its variants Refa-InfoNCA (dynamic) and Refa-InfoNCA (1-vs-all) to assess the impact of dynamic adaptation and 1-vs-all contrastive optimization, as summarized in Table 1. The results demonstrate that Refa-InfoNCA (1-vs-all) outperforms both Refa-InfoNCA (dynamic) and the baseline Refa-InfoNCA across all metrics. Specifically, Refa-InfoNCA (1-vs-all) achieves slight improvement (nearly 1%) over Refa-InfoNCA (dynamic) and 2.7% on LC-WR and 3.9% on WR over the baseline. These improvements validate the theoretical advantages of the 1-vs-all contrastive optimization framework, which ensures better utilization of negative samples by contrasting each positive instance against all available negatives, leading to more robust preference alignment. Furthermore, the superior performance of Refa-InfoNCA (dynamic) over the baseline underscores the effectiveness of incorporating instance-specific dynamic adjustments. Our theoretical in (Appendix D) and empirical analysis corroborates these findings by proving that the 1-vs-all strategy enhances the representational capacity of the model, resulting in superior generalization and preference alignment compared to other baselines.
Refa-1-vs-all vs Refa-dynamic
We compared the performance of Refa-1-vs-all and Refa-dynamic, as shown in Table 1, to evaluate the impact of dynamic adaptation in preference optimization. Refa-dynamic outperforms Refa-1-vs-all across all metrics, achieving improvement of 2.9% (LC-WR) and 3.4% (WR) over Refa-1-vs-all. These results highlight the superiority of Refa-dynamic method, which dynamically adjusts preference weights and responses to improve alignment. Theoretical analysis in (Appendix E) further substantiates REFA-Dynamic to adapt more flexibly to diverse input contexts, resulting in enhanced generalization and preference alignment compared to the static 1-vs-all framework.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Supplementary Materials
  These supplementary materials provide additional details, derivations, and experimental results for our paper. The appendix is organized as follows:
  • Section A provides a detailed overview of related work pertaining to this paper.
  • Section B analyzes the InfoNCA loss function, deriving its gradients, characterizing its stationary points, and examining the role of the reference model.
  • Section C provides additional analysis of the InfoNCA loss in a reference-free setting, highlighting issues like promoting low-quality responses and generating contradictory optimization signals.
  • Section D details how addressing identified weaknesses in the reference-free InfoNCA formulation leads to a refined, reference-free multi-preference loss function, paving the way for the proposed Refa-base variants.
  • Section E introduces the Refa-base loss functions without length normalization, explaining how they incorporate multiple preferences and reward-based weighting to overcome prior limitations.
  • Section F focuses on length normalization in Refa, demonstrating how normalizing per-token probabilities prevents trivial short-response biases and encourages genuinely improved responses.
  • Section G explores the nuances in the relationship between average per-token uncertainty and sequence length, and justifies the addition of an EOS-probability regularizer to counter subtle length-related incentives.
  • Section H More detailed discussion about the Baselines and Evaluation Benchmarks.
Appendix A Related Work
A.1 Preference Optimization for Alignment.
Early alignment efforts focussed on RLHF through learning an intermediate reward model using reinforcement learning algorithms like PPO (Schulman et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). While effective, this can be computationally expensive and noisy due to the intermediate reward estimation. Recent approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), streamline alignment by directly optimizing a contrastive loss over pairs of preferred and dispreferred responses, bypassing explicit reward models. Subsequent works have extended this idea, exploring variants like IPO (Azar et al., 2023), CPO (Xu et al., 2024a), ORPO (Hong et al., 2024a), and R-DPO (Park et al., 2024), each offering alternative formulations or regularizations. Additionally, methods like RRHF (Yuan et al., 2023), SLiC-HF (Zhao et al., 2023a), KTO (Ethayarajh et al., 2024), and SimPO (Meng et al., 2024) propose diverse preference optimization objectives, some relying on reference models and others operating reference-free. A recent unifying framework by Tang et al. (2024) shows that many such offline algorithms can be viewed as special cases of a generalized preference optimization approach.
A.2 Reference-Free Alignment.
While many alignment techniques rely on a reference model (Ouyang et al., 2022), there is growing interest in reference-free methods that directly optimize from scalar feedback. Reference-free approaches avoid the complexity of ratio modeling and can leverage richer data containing multiple candidate responses per query (Cui et al., 2023; Liu et al., 2024; Zhou et al., 2023; Wang et al., 2024). Recently, objectives like SimPO (Meng et al., 2024) have shown that focusing on the log-probability of sequences without a separate reference model can achieve strong performance, making the training pipeline simpler and potentially more robust.
A.3 Multi-Preference Optimization.
The development of multi-preference and reference-free approaches is facilitated by datasets like the UltraFeedback dataset (Cui et al., 2023) that provide scalar rewards corresponding to multiple responses related to a single query. Methods like InfoNCA (Chen et al., 2024) and SWEPO (Gupta et al., 2024) use such datasets to generalize beyond pairwise comparisons, simultaneously considering sets of positive and negative responses. Multi-preference objectives better approximate the true preference landscape, better enabling models to estimate the accepted and rejected distributions. While InfoNCA leverages a noise-contrastive framework to align responses according to scalar rewards, SWEPO introduces deviation-based weighting to emphasize highly positive or highly negative deviations more strongly, thus improving alignment quality.
In summary, our work extends the literature on reference-free, multi-preference optimization methods. By integrating length normalization, deviation-based weighting, and EOS probability regularization, we build on principles introduced by methods like InfoNCA and SWEPO to address subtle length-based incentives, ultimately producing richer and more human-aligned responses.
Appendix B Analysis of the InfoNCA Loss
In this section, we provide a detailed analysis of the InfoNCA loss function introduced in previous works. We begin by establishing our notation and problem setup, then proceed to define the InfoNCA loss. Subsequently, we derive its gradients, characterize its stationary points, and examine the implications for the learned policy distribution when a reference distribution is present. Finally, we discuss how removing the reference model recovers a simpler scenario that aligns the learned distribution directly with the target distribution.
B.1 Notation and Setup
We consider a query (or context) $x\in\mathcal{X}$ and a set of $K$ candidate responses $\{y_{i}\}_{i=1}^{K}$. Each response $y_{i}$ is associated with a scalar reward $r_{i}=r(x,y_{i})$, representing how suitable or aligned the response is according to some evaluation metric or annotated feedback.
We assume access to:
  • A policy model $\pi_{\theta}(y|x)$ parameterized by $\theta$, which assigns a probability to each response $y_{i}$.
  • A reference model $\mu(y|x)$, which provides a baseline distribution over responses. This reference model is fixed and not optimized.
We define the ratio:
$r_{\theta}(x,y_{i})\;:=\;\log\frac{\pi_{\theta}(y_{i}|x)}{\mu(y_{i}|x)}.$ (15)
For simplicity, we set $\beta=1$ in the definitions that follow, but the analysis easily generalizes to arbitrary positive $\beta$.
The target distribution derived from the rewards is given by a softmax transformation:
$p_{i}^{\text{target}}\;:=\;\frac{e^{\alpha r_{i}}}{\sum_{j=1}^{K}e^{\alpha r_{j}}},$ (16)
where $\alpha>0$ is an inverse temperature-like parameter controlling the sharpness of the target distribution.
B.2 The InfoNCA Loss
The InfoNCA loss aims to match the distribution induced by $r_{\theta}(x,y)$ to the target distribution $p_{i}^{\text{target}}$. Given that $r_{\theta}(x,y_{i})=\log(\pi_{\theta}(y_{i}|x)/\mu(y_{i}|x))$, we define the model distribution as:
$p_{i}^{\text{model}}\;:=\;\frac{e^{r_{\theta}(x,y_{i})}}{\sum_{j=1}^{K}e^{r_{\theta}(x,y_{j})}}.$ (17)
The InfoNCA loss is the cross-entropy between the target distribution $p_{i}^{\text{target}}$ and the model distribution $p_{i}^{\text{model}}$:
$L_{\text{InfoNCA}}(\theta)\;=\;-\sum_{i=1}^{K}p_{i}^{\text{target}}\,\log p_{i}^{\text{model}}.$ (18)
Minimizing $L_{\text{InfoNCA}}$ encourages $p_{i}^{\text{model}}$ to align with $p_{i}^{\text{target}}$.
B.3 Gradient Derivation
To understand the optimization dynamics, we compute the gradient of $L_{\text{InfoNCA}}$ with respect to the log-ratios $r_{\theta}(x,y_{i})$.
First, note that:
$\log p_{i}^{\text{model}}\;=\;r_{\theta}(x,y_{i})-\log\biggl{(}\sum_{j=1}^{K}e^{r_{\theta}(x,y_{j})}\biggr{)}.$
Differentiating $\log p_{k}^{\text{model}}$ with respect to $r_{\theta}(x,y_{i})$:
$\frac{\partial\log p_{k}^{\text{model}}}{\partial r_{\theta}(x,y_{i})}\;=\;\delta_{ik}-p_{i}^{\text{model}},$
where $\delta_{ik}=1$ if $i=k$ and 00 otherwise.
Now, differentiate the loss:
$\frac{\partial L_{\text{InfoNCA}}}{\partial r_{\theta}(x,y_{i})}\;=\;-\sum_{k=1}^{K}p_{k}^{\text{target}}\frac{\partial\log p_{k}^{\text{model}}}{\partial r_{\theta}(x,y_{i})}.$
Substituting the derivative of $\log p_{k}^{\text{model}}$:
$\frac{\partial L_{\text{InfoNCA}}}{\partial r_{\theta}(x,y_{i})}=-\sum_{k=1}^{K}p_{k}^{\text{target}}(\delta_{ik}-p_{i}^{\text{model}}).$
Evaluating the sum yields the well-known cross-entropy gradient form:
$\frac{\partial L_{\text{InfoNCA}}}{\partial r_{\theta}(x,y_{i})}\;=\;p_{i}^{\text{model}}-p_{i}^{\text{target}}.$
Thus, the gradient simplifies to the difference between the model and target distributions.
B.4 Stationary Points
A stationary point occurs where the gradient is zero for all $i$:
$p_{i}^{\text{model}}-p_{i}^{\text{target}}=0\quad\implies\quad p_{i}^{\text{model}}=p_{i}^{\text{target}}\;\;\forall i.$
At this stationary point, the ratio distribution $e^{r_{\theta}(x,y)}/\sum_{j}e^{r_{\theta}(x,y_{j})}$ perfectly matches the reward-derived target distribution. However, note that $p_{i}^{\text{model}}$ is defined on the ratio $\pi_{\theta}(y|x)/\mu(y|x)$, not directly on $\pi_{\theta}(y|x)$ alone.
Implications for the Ratio $\pi_{\theta}(y|x)/\mu(y|x)$
Recall:
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)/\mu(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)/\mu(y_{j}|x)}.$
At the stationary point:
$p_{i}^{\text{target}}=p_{i}^{\text{model}}=\frac{\frac{\pi_{\theta}(y_{i}|x)}{\mu(y_{i}|x)}}{\sum_{j=1}^{K}\frac{\pi_{\theta}(y_{j}|x)}{\mu(y_{j}|x)}}.$
Rearranging:
$\frac{\pi_{\theta}(y_{i}|x)}{\mu(y_{i}|x)}=p_{i}^{\text{target}}\sum_{j=1}^{K}\frac{\pi_{\theta}(y_{j}|x)}{\mu(y_{j}|x)}.$
Define:
$A:=\sum_{j=1}^{K}\frac{\pi_{\theta}(y_{j}|x)}{\mu(y_{j}|x)}.$
Then:
$\pi_{\theta}(y_{i}|x)=p_{i}^{\text{target}}\mu(y_{i}|x)A.$
This shows that at equilibrium, $\pi_{\theta}(y_{i}|x)$ is proportional to the product $p_{i}^{\text{target}}\mu(y_{i}|x)$. Note that we have not assumed $\sum_{i}\pi_{\theta}(y_{i}|x)=1$; $\pi_{\theta}(y_{i}|x)$ could be any nonnegative values. The shape of the solution is defined by:
$\pi_{\theta}(y_{i}|x)\propto p_{i}^{\text{target}}\mu(y_{i}|x).$
If we subsequently impose normalization to interpret $\pi_{\theta}(y_{i}|x)$ as a probability distribution, we would set:
$\hat{\pi}_{\theta}(y_{i}|x)=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)}=\frac{p_{i}^{\text{target}}\mu(y_{i}|x)}{\sum_{k=1}^{K}p_{k}^{\text{target}}\mu(y_{k}|x)}.$
Define:
$M:=\sum_{k=1}^{K}p_{k}^{\text{target}}\mu(y_{k}|x),$
then:
$\hat{\pi}_{\theta}(y_{i}|x)=\frac{p_{i}^{\text{target}}\mu(y_{i}|x)}{M}.$
In this normalized view, the equilibrium policy distribution aligns with a $\mu$-weighted version of the target distribution rather than $p_{i}^{\text{target}}$ directly. This may point to a potential improvement over the InfoNCA loss function, simply by removing the reference model, as we show in the subsection below.
B.5 Removing the Reference Model: A Reference-Free Scenario
The presence of the reference model $\mu(y|x)$ skews the stationary solution. If we were to remove $\mu(y|x)$ from the formulation—i.e., consider a scenario where $r_{\theta}(x,y):=\log\pi_{\theta}(y|x)$ without dividing by $\mu(y|x)$—then the model distribution $p_{i}^{\text{model}}$ would directly be:
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)},$
which is simply the policy distribution itself.
In that reference-free case, setting $p_{i}^{\text{model}}=p_{i}^{\text{target}}$ at the stationary point would imply:
$\pi_{\theta}(y_{i}|x)=p_{i}^{\text{target}},$
directly aligning the policy distribution with the target distribution. This is the desirable outcome if no baseline or reference distribution is involved. Thus, the introduction of $\mu(y|x)$ in the denominator shifts the stationary solution, necessitating careful design if we want direct alignment with $p_{i}^{\text{target}}$.
In summary, the InfoNCA loss aligns the ratio $\pi_{\theta}(y|x)/\mu(y|x)$ to the target $p_{i}^{\text{target}}$. To achieve direct alignment of $\pi_{\theta}(y|x)$ with $p_{i}^{\text{target}}$, a reference-free formulation is required.
Appendix C Additional Analysis of the InfoNCA Loss Without a Reference Model
In the previous section, we analyzed the InfoNCA loss in the presence of a reference model $\mu(y|x)$. Here, we revisit the InfoNCA formulation under a simpler, reference-free scenario and highlight several issues that arise when optimizing this objective. Specifically, we show how the loss may inadvertently increase the probabilities of low-rated responses and lead to contradictory optimization signals.
C.1 Reference-Free InfoNCA
Removing the reference model from the formulation amounts to setting:
$r_{\theta}(x,y_{i})\;:=\;\log\pi_{\theta}(y_{i}|x).$ (19)
In this case, the model distribution becomes:
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}.$ (20)
The InfoNCA loss reduces to the standard cross-entropy between the target and model distributions:
$L_{\text{InfoNCA}}(\theta)\;=\;-\sum_{i=1}^{K}p_{i}^{\text{target}}\log p_{i}^{\text{model}},$ (21)
with $p_{i}^{\text{target}}$ defined as before.
C.2 Term-by-Term Gradient Dynamics
Lemma C.1 (Gradient of a Single-Term Objective).
Let $\ell_{i}(\theta)$ be defined as:
$\ell_{i}(\theta):=-p_{i}^{\text{target}}\log p_{i}^{\text{model}},$
where
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}$
and $p_{i}^{\text{target}}$ is a fixed scalar satisfying $p_{i}^{\text{target}}>0$. Assume $\pi_{\theta}(y_{j}|x)>0$ for all $j=1,\dots,K$.
Then, the partial derivatives of $\ell_{i}(\theta)$ with respect to $\pi_{\theta}(y_{i}|x)$ and $\pi_{\theta}(y_{j}|x)$ for $j\neq i$ are:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}=-p_{i}^{\text{target}}\left(\frac{1}{\pi_{\theta}(y_{i}|x)}-\frac{1}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}\right),$
and for each $j\neq i$:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}=p_{i}^{\text{target}}\frac{1}{\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)}.$
Proof.
We start from the definition:
$\ell_{i}(\theta)=-p_{i}^{\text{target}}\log p_{i}^{\text{model}}.$
Substitute $p_{i}^{\text{model}}$:
$p_{i}^{\text{model}}=\frac{\pi_{\theta}(y_{i}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}.$
Thus:
$\ell_{i}(\theta)=-p_{i}^{\text{target}}\left[\log\pi_{\theta}(y_{i}|x)-\log\left(\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)\right)\right].$
First, consider the derivative with respect to $\pi_{\theta}(y_{i}|x)$:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}=-p_{i}^{\text{target}}\left[\frac{\partial}{\partial\pi_{\theta}(y_{i}|x)}\log\pi_{\theta}(y_{i}|x)-\frac{\partial}{\partial\pi_{\theta}(y_{i}|x)}\log\left(\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)\right)\right].$
Compute each derivative separately:
1. For the first term:
$\frac{\partial}{\partial\pi_{\theta}(y_{i}|x)}\log\pi_{\theta}(y_{i}|x)=\frac{1}{\pi_{\theta}(y_{i}|x)}.$
2. For the second term, let $S:=\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)$. Then:
$\frac{\partial}{\partial\pi_{\theta}(y_{i}|x)}\log(S)=\frac{1}{S}\frac{\partial S}{\partial\pi_{\theta}(y_{i}|x)}=\frac{1}{S}.$
Substituting back:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}=-p_{i}^{\text{target}}\left(\frac{1}{\pi_{\theta}(y_{i}|x)}-\frac{1}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}\right).$
Next, consider $j\neq i$:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}=-p_{i}^{\text{target}}\left[0-\frac{\partial}{\partial\pi_{\theta}(y_{j}|x)}\log\left(\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)\right)\right].$
Since $\log(S)$ with $S=\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)$ gives:
$\frac{\partial}{\partial\pi_{\theta}(y_{j}|x)}\log(S)=\frac{1}{S}.$
Therefore:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}=p_{i}^{\text{target}}\frac{1}{\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)}.$
This completes the proof. ∎
Lemma C.2 (Directional Influence of a Single-Term Objective).
Under the same assumptions as Lemma C.1, consider the single-term objective $\ell_{i}(\theta)$ defined above. To reduce $\ell_{i}(\theta)$, one should:
  • Increase $\pi_{\theta}(y_{i}|x)$: The partial derivative $\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}$ can be negative if $\pi_{\theta}(y_{i}|x)$ is sufficiently small, indicating that increasing $\pi_{\theta}(y_{i}|x)$ will decrease $\ell_{i}(\theta)$.
  • Decrease $\pi_{\theta}(y_{j}|x)$ for $j\neq i$: For each $j\neq i$, the partial derivative $\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}$ is strictly positive. Thus, increasing $\pi_{\theta}(y_{j}|x)$ raises $\ell_{i}(\theta)$, while decreasing it lowers $\ell_{i}(\theta)$.
In summary, minimizing $\ell_{i}(\theta)$ encourages increasing the probability of the $i$-th response and reducing the probabilities of all other responses.
Proof.
From Lemma C.1, we have:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{i}|x)}=-p_{i}^{\text{target}}\left(\frac{1}{\pi_{\theta}(y_{i}|x)}-\frac{1}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)}\right).$
Since $p_{i}^{\text{target}}>0$, the sign of this derivative depends on the relative sizes of $\pi_{\theta}(y_{i}|x)$ and $\sum_{j}\pi_{\theta}(y_{j}|x)$. If $\pi_{\theta}(y_{i}|x)$ is small compared to the total sum, then:
$\frac{1}{\pi_{\theta}(y_{i}|x)}>\frac{1}{\sum_{j}\pi_{\theta}(y_{j}|x)},$
making the entire expression negative. A negative derivative w.r.t. $\pi_{\theta}(y_{i}|x)$ implies increasing $\pi_{\theta}(y_{i}|x)$ reduces $\ell_{i}(\theta)$.
On the other hand, for $j\neq i$:
$\frac{\partial\ell_{i}(\theta)}{\partial\pi_{\theta}(y_{j}|x)}=p_{i}^{\text{target}}\frac{1}{\sum_{k=1}^{K}\pi_{\theta}(y_{k}|x)}>0.$
Since $p_{i}^{\text{target}}>0$ and the denominator is positive, this derivative is strictly positive. Thus, increasing $\pi_{\theta}(y_{j}|x)$ with $j\neq i$ increases $\ell_{i}(\theta)$, and decreasing $\pi_{\theta}(y_{j}|x)$ reduces $\ell_{i}(\theta)$.
Combining these observations, to minimize $\ell_{i}(\theta)$, we must increase $\pi_{\theta}(y_{i}|x)$ (when beneficial) and decrease $\pi_{\theta}(y_{j}|x)$ for all $j\neq i$. This establishes the stated directional influence on the single-term objective. ∎
Appendix D Fixing Weaknesses in InfoNCA Style Loss Leads to Reference Free Swepo
In the absence of a reference model, the InfoNCA loss reduces to a standard cross-entropy form that aligns the model distribution with a target distribution derived from rewards. Although seemingly straightforward, this formulation exhibits several critical weaknesses when applied to multiple responses with varying quality. In this section, we highlight these issues step-by-step, referencing Lemma C.1 and Lemma C.2, which describe how each individual loss term behaves, and propose incremental fixes that ultimately motivate a more refined approach.
D.1 W1: Encouraging Low-Rated Responses
Issue: From Lemma C.1 and Lemma C.2, each single-term objective $\ell_{i}(\theta)$ encourages increasing $\pi_{\theta}(y_{i}|x)$ regardless of the response’s quality. Even if $p_{i}^{\text{target}}$ is small, as long as it is nonzero, the model is pushed to raise the probability of that response, effectively giving low-rated responses unwarranted attention.
Implication: Low-quality responses should not receive reinforcement, yet the InfoNCA-style objective provides a positive gradient push as long as $p_{i}^{\text{target}}>0$. This introduces noise into training and can skew the model toward suboptimal responses.
Fix 1: Removing Terms for Low-Rated Responses
A natural remedy is to remove terms corresponding to low-reward responses. For instance, define a threshold (e.g., mean reward $\bar{r}$) and restrict the loss to responses above this threshold:
$\widetilde{L}(\theta):=-\sum_{i\in\mathcal{I}_{\text{pos}}}p_{i}^{\text{target}}\log p_{i}^{\text{model}},\quad\mathcal{I}_{\text{pos}}=\{i\mid r_{i}>\bar{r}\}.$
Excluding low-rated responses avoids their direct encouragement. However, this approach introduces another weakness.
D.2 W2: Contradictory Signals Among Good Responses
Issue: After removing low-rated responses, only “good” responses remain. Yet, Lemma C.2 still applies: each remaining term $\ell_{i}(\theta)$ attempts to boost its own response’s probability while suppressing all others, including other good ones. Consider responses $\{8,7,3,2\}$ and exclude those below the mean. The terms for $8$-rated and $7$-rated responses now compete against each other, creating contradictory optimization signals.
Implication: We want multiple good responses to coexist and improve together. However, the pairwise competitive nature of these single-term objectives prevents a collective uplift. Instead of reinforcing a set of good responses, the model ends up with internal conflicts, slowing training and potentially leading to suboptimal solutions.
Fix 2: Single “1 vs All” Term
One idea is to avoid summing over multiple per-response objectives. Instead, choose one top-rated response and form a “1 vs all” loss:
$L_{\text{1vsAll}}(\theta)=-\log\frac{\pi_{\theta}(y_{i^{*}}|x)}{\sum_{j=1}^{K}\pi_{\theta}(y_{j}|x)},$
where $y_{i^{*}}$ is the best response. This removes direct conflicts among multiple positive responses since we focus on only one. But what if there are multiple top-tier responses we wish to jointly learn from?
D.3 W3: Handling Multiple Equally Good Responses is Not Possible with 1 vs All
Issue: If there are multiple equally good responses, say $\{8,8,1,1\}$, the “1 vs all” approach cannot leverage both top responses simultaneously. It only promotes a single chosen response $y_{i^{*}}$ at the expense of all others. If we pick the first “8”-rated response, we lose the opportunity to learn from the second “8”-rated one.
Implication: The “1 vs all” fix addresses contradictory signals when multiple good responses are present only by ignoring some of them, which is not ideal. We want a mechanism that can consider all top-tier responses together, promoting them as a group.
Fix 3: Swepo-Style Single Term (All Good in Numerator, All in Denominator)
We can construct a single-term loss function that includes all positively rated responses in the numerator and all responses (both positive and negative) in the denominator. This resembles a swepo-like contrastive form:
$L_{\text{all-pos}}(\theta)=-\log\frac{\sum_{y\in Y^{+}}\pi_{\theta}(y|x)}{\sum_{y\in Y^{+}\cup Y^{-}}\pi_{\theta}(y|x)},$
where $Y^{+}$ are all good responses and $Y^{-}$ are all non-positive responses. This addresses the coexistence issue, but now every positive and negative is treated equally, ignoring the magnitude of deviations.
D.4 W4: Lack of Weighting
Issue: In the swepo-style single-term approach without weighting, all positive responses are treated identically, as are all negative responses. Consider ratings $\{10,4,4,1\}$: $1$ is significantly worse than $4$, yet both are just “negative” if below the mean threshold. In other cases, like $\{9,7,3,1\}$, $9$ and $7$ are weighted equally, without learning more from $9$ than $7$. Without weighting, we fail to differentiate these nuanced gradations in quality.
Implication: We want to reward the most outstanding positive responses more than marginally positive ones. Similarly, the truly poor responses should have a stronger negative influence than those that are just below average. Ignoring this leads to suboptimal gradient signals that do not fully exploit the granularity of the reward information.
Fix 4: Deviation-Based Weighting
Assign weights based on the deviation from the mean reward. Responses with higher (positive) deviation get amplified; those with negative deviation are penalized proportionally. This yields:
$L_{\text{weighted}}(\theta)=-\log\frac{\sum_{y\in Y^{+}}w_{y}\pi_{\theta}(y|x)}{\sum_{y\in Y^{+}\cup Y^{-}}w_{y}\pi_{\theta}(y|x)},$
where $w_{y}=e^{\alpha\Delta S_{y}}$ where $\Delta S_{y}=(r_{y}-\overline{r})^{p}$ for some p-value like $p=\{1,2\}$, where $\overline{r}$ is the mean reward. This weighting refines the loss by making the gradient more informative and aligned with the underlying quality differences.
D.5 W5: Fine-Grained Control via Hyperparameters
Issue: While deviation-based weighting improves the distribution of gradient signals, we may still need more control. Different tasks might require adjusting how heavily to penalize negatives, how sharply to differentiate based on deviations, or how to set an effective “temperature” of the softmax.
Implication: A flexible loss function should allow fine-tuning how negatives compete with positives ($\gamma$ parameter), how sensitive it is to rating differences ($\alpha$ parameter), and how sharply it distinguishes between responses ($\beta$ parameter).
Fix 5: Hyperparameterized Weighted Multi-Preference Loss
By introducing hyperparameters $(\alpha,\beta,\gamma)$, we can regulate:
  1. 1.
    $\alpha$: how strongly the deviation affects weighting.
  2. 2.
    $\beta$: an inverse-temperature parameter controlling the softmax sharpness.
  3. 3.
    $\gamma$: a factor that overweights negative responses to ensure a clear contrast.
Incorporating these parameters, we obtain a final, fully adjustable loss function that addresses all previously identified weaknesses, allowing nuanced and stable optimization.
$L_{\text{swepo}}(\theta)=-\log\frac{\sum_{y\in Y^{+}}e^{\beta(\pi_{\theta}(y|x)+\alpha\Delta S_{y})}}{\sum_{y\in Y^{+}}e^{\beta(\pi_{\theta}(y|x)+\alpha\Delta S_{y})}+\gamma\sum_{y\in Y^{-}}e^{\beta(\pi_{\theta}(y|x)+\alpha\Delta S_{y})}}$
In summary, each weakness of the InfoNCA-style loss reveals the necessity of more sophisticated mechanisms: filtering out low-rated responses, avoiding contradictory objectives among top responses, considering multiple good responses together, weighting by deviation for finer quality distinctions, and introducing hyperparameters for greater flexibility. These insights set the stage for refined multi-preference optimization methods that align more closely with the complexity of real-world preference alignment scenarios.
Appendix E Proposed RefaBase Loss Functions Without Length Normalization
The analysis in the previous sections uncovers several fundamental weaknesses in an InfoNCA-style loss when used without a reference model. We identified:
  1. 1.
    Encouragement of low-rated responses (W1): Every response, including low-quality ones, receives upward pressure as long as $p_{i}^{\text{target}}>0$, leading to suboptimal reinforcement.
  2. 2.
    Conflicts among top-tier responses (W2 & W3): Even after filtering out low-quality responses, multiple good responses compete against each other. This prevents a collective uplift and complicates the optimization dynamics.
  3. 3.
    Lack of nuanced weighting (W4): Treating all positive or negative responses equally ignores the fine-grained quality differences indicated by their rewards, missing opportunities to better leverage the available information.
  4. 4.
    Insufficient flexibility (W5): Without hyperparameters, one cannot adjust how strongly to penalize negatives, how sensitively to incorporate deviations, or how sharply to distinguish among responses.
These issues highlight the need for a more refined reference-free multi-preference loss function that:
  1. 1.
    Prioritizes genuinely high-quality responses without inadvertently promoting low-quality ones.
  2. 2.
    Reduces direct competition among good responses so that multiple top-tier candidates can improve collectively.
  3. 3.
    Incorporates reward-based weighting to reflect the spectrum of response quality.
  4. 4.
    Provides hyperparameter-driven flexibility for controlling emphasis on negatives, sensitivity to rating differences, and the sharpness of the distribution.
E.1 RefaBase Loss Variants
We propose a family of RefaBase (Reference-Free Alignment) loss functions that build upon these insights. By integrating deviation-based weighting, selective response sets, and hyperparameters for fine-grained control, we address the aforementioned weaknesses. We present two variants:
RefaBase-1 vs All:
This variant selects the single highest-rated response $y_{i^{*}}$ defined as:
$y_{i^{*}}=\arg\max_{i}r_{i},\quad Y^{+}=\{y_{i^{*}}\},\quad Y^{-}=\{y_{j}\mid j\neq i^{*}\}.$
We assign weights based on deviation from the mean, $\Delta S_{i}=r_{i}-\bar{r}$, using an exponential scheme $w_{i}=e^{\alpha\Delta S_{i}}$. Introducing $\beta$ for inverse-temperature and $\gamma$ for negative overweighting, the RefaBase-1 vs All loss is:
[…]
This approach eliminates low-rated responses and avoids direct conflict among multiple good responses by focusing solely on the top candidate. However, it cannot simultaneously exploit multiple equally high-quality responses.
RefaBase-Dynamic:
To incorporate multiple good responses, we define:
$Y^{+}=\{y_{i}\mid r_{i}>\bar{r}\},\quad Y^{-}=\{y_{j}\mid r_{j}\leq\bar{r}\}.$
All positive responses appear in the numerator, and both positive and negative sets appear in the denominator. Applying deviation-based weights and hyperparameters as before, we obtain a swepo-style loss:
[…]
This RefaBase-dynamic formulation simultaneously addresses all previously identified weaknesses:
  • W1: Low-rated responses remain in the denominator but do not receive direct promotion.
  • W2 & W3: Multiple top-tier responses are collectively improved, reducing pairwise conflicts.
  • W4: Deviation-based weighting differentiates among positive responses and among negative ones.
  • W5: Hyperparameters $(\alpha,\beta,\gamma)$ enable nuanced control over the optimization process.
In essence, the RefaBase-dynamic loss represents the final convergence of our incremental fixes, providing a robust and flexible solution to the shortcomings of the InfoNCA-style loss in reference-free, multi-preference alignment scenarios.
Appendix F Length Normalization in RefaBase
In this section, we analyze the behavior of a simplified RefaBase loss function without hyperparameters or weighting. We focus on the dynamic variant, which selects positive and negative responses based on their rewards relative to the mean. We then highlight how length can become a shortcut for reducing the loss and present a normalization strategy to address this issue.
F.1 Simplified RefaBase-Dynamic Loss
Consider the simplified RefaBase-dynamic loss:
[…]
where
  • $Y^{+}=\{y\mid r_{y}>\bar{r}\}$ is the set of responses with rewards above the mean $\bar{r}$.
  • $Y^{-}=\{y\mid r_{y}\leq\bar{r}\}$ is the set of responses with rewards below or equal to the mean.
  • $\pi_{\theta}(y|x)=e^{s_{y}}$, where $s_{y}=\log\pi_{\theta}(y|x)$ is the log-probability of response $y$.
Define:
$Z^{+}:=\sum_{y\in Y^{+}}\pi_{\theta}(y|x),\quad Z:=\sum_{y\in Y^{+}\cup Y^{-}}\pi_{\theta}(y|x).$
Thus:
[…]
F.2 Gradient Analysis of the Simplified RefaBase-Dynamic Loss
Lemma F.1 (Gradient of the Simplified RefaBase-Dynamic Loss).
[…]
where
$p_{y}^{\text{model}}=\frac{\pi_{\theta}(y|x)}{Z},\quad\text{and}\quad p_{y}^{\text{pos}}=\frac{\pi_{\theta}(y|x)}{Z^{+}}.$
Proof.
First, note:
[…]
with
$Z=\sum_{y\in Y^{+}}\pi_{\theta}(y|x)+\sum_{y\in Y^{-}}\pi_{\theta}(y|x),\quad Z^{+}=\sum_{y\in Y^{+}}\pi_{\theta}(y|x).$
Taking derivatives w.r.t. $s_{y}$:
$\frac{\partial L}{\partial s_{y}}=\frac{\partial\log Z}{\partial s_{y}}-\frac{\partial\log Z^{+}}{\partial s_{y}}.$
Compute each term:
$\frac{\partial\log Z}{\partial s_{y}}=\frac{1}{Z}\frac{\partial Z}{\partial s_{y}},\quad\frac{\partial\log Z^{+}}{\partial s_{y}}=\frac{1}{Z^{+}}\frac{\partial Z^{+}}{\partial s_{y}}.$
If $y\in Y^{+}$:
$\frac{\partial Z^{+}}{\partial s_{y}}=\pi_{\theta}(y|x)=e^{s_{y}},\quad\frac{\partial Z}{\partial s_{y}}=\pi_{\theta}(y|x).$
Thus:
$\frac{\partial L}{\partial s_{y}}=\frac{\pi_{\theta}(y|x)}{Z}-\frac{\pi_{\theta}(y|x)}{Z^{+}}=p_{y}^{\text{model}}-p_{y}^{\text{pos}}.$
If $y\in Y^{-}$:
$\frac{\partial Z^{+}}{\partial s_{y}}=0,\quad\frac{\partial Z}{\partial s_{y}}=\pi_{\theta}(y|x).$
Thus:
$\frac{\partial L}{\partial s_{y}}=\frac{\pi_{\theta}(y|x)}{Z}-0=p_{y}^{\text{model}}.$
This completes the proof. ∎
F.3 Implications for Positive and Negative Responses
Lemma F.2 (Increasing Probability of Positive Responses Decreases the Loss).
[…]
Proof.
Recall the simplified RefaBase-dynamic loss:
[…]
where
$Z^{+}=\sum_{y\in Y^{+}}\pi_{\theta}(y|x),\quad Z=Z^{+}+\sum_{y\in Y^{-}}\pi_{\theta}(y|x).$
Since $Y^{+}\subseteq Y^{+}\cup Y^{-}$, it follows that $Z^{+}\leq Z$, and strictly $Z^{+}<Z$ if there is at least one negative response.
Case $y\in Y^{+}$:
Notice that:
$p_{y}^{\text{model}}:=\frac{\pi_{\theta}(y|x)}{Z},\quad p_{y}^{\text{pos}}:=\frac{\pi_{\theta}(y|x)}{Z^{+}}.$
Since $Z^{+}<Z$, we have $\frac{1}{Z^{+}}>\frac{1}{Z}$, implying $p_{y}^{\text{pos}}>p_{y}^{\text{model}}$. Therefore:
$\frac{\partial L}{\partial s_{y}}=p_{y}^{\text{model}}-p_{y}^{\text{pos}}<0.$
A negative gradient w.r.t. $s_{y}$ means that increasing $s_{y}$ (equivalently, increasing $\pi_{\theta}(y|x)$) for $y\in Y^{+}$ decreases the loss $L$. In other words, raising the probability of a positive response lowers the loss.
Case $y\in Y^{-}$:
If $y\in Y^{-}$, it does not appear in $Z^{+}$. Thus:
$\frac{\partial Z^{+}}{\partial s_{y}}=0,\quad\frac{\partial Z}{\partial s_{y}}=\pi_{\theta}(y|x).$
Then:
$\frac{\partial L}{\partial s_{y}}=\frac{\pi_{\theta}(y|x)}{Z}=p_{y}^{\text{model}}>0.$
A positive gradient w.r.t. $s_{y}$ indicates that increasing $s_{y}$ (or $\pi_{\theta}(y|x)$) for a negative response $y$ raises the loss $L$. To reduce the loss, we must not increase the probabilities of negative responses.
Conclusion:
[…]
F.4 Length as a Shortcut
Without length normalization, $\log\pi_{\theta}(y|x)$ is the sum of token log-probabilities:
$\log\pi_{\theta}(y|x)=\sum_{t\in y}\log P_{\theta}(t|\text{context}).$
Shorter responses generally have fewer tokens over which to accumulate negative log-probabilities, often resulting in higher $\pi_{\theta}(y|x)$. Thus, a trivial way to increase $\pi_{\theta}(y|x)$ for some $y\in Y^{+}$ is to make $y$ exceedingly short, increasing its probability without genuinely improving per-token quality.
Lemma F.3 (Length-Induced Probability Inflation).
[…]
Proof.
If each token in both responses has an expected log-probability around $-c$ for some $c>0$, then:
$\log\pi_{\theta}(y_{1}|x)\approx-|y_{1}|c,\quad\log\pi_{\theta}(y_{2}|x)\approx-|y_{2}|c.$
If $|y_{1}|<|y_{2}|$, then $-|y_{1}|c>-|y_{2}|c$ and thus:
$\pi_{\theta}(y_{1}|x)=e^{-|y_{1}|c}>e^{-|y_{2}|c}=\pi_{\theta}(y_{2}|x).$
This shows shorter responses achieve higher probabilities, providing an artificial advantage. ∎
F.5 Fix: Length Normalization
To prevent the model from exploiting this length-based shortcut, we introduce length normalization. Instead of:
$\log\pi_{\theta}(y|x)=\sum_{t\in y}\log P_{\theta}(t|\text{context}),$
we define:
$\overline{\log\pi_{\theta}}(y|x)=\frac{1}{|y|}\sum_{t\in y}\log P_{\theta}(t|\text{context}),$
and use $e^{\overline{\log\pi_{\theta}}(y|x)}$ in place of $\pi_{\theta}(y|x)$ in the loss. This ensures that shorter responses do not gain an undue advantage from having fewer tokens, compelling the model to improve the quality per token rather than simply reducing response length.
F.6 Modified Loss Function
Given the above length normalization fix, we can now modify our loss function to use the new length normalized log-perplexity values, rather than simply the log-perplexities. We write this as follows:
[…]
Appendix G Relationship between Average Uncertainty and Sequence Length
Despite our length normalization, we find that the loss function given by 26 still has problems due to a nuanced observation. We will state this observation here as a conjecture, and verify it through our experiments. But before we state this observation, we provide some notations.
Notation: Let $y$ denote a response (sequence of tokens) with length $\operatorname{len}(y)$ and token-level conditional probabilities $P_{\theta}(t\mid\text{context})$. We define the average per-token negative log-probability (perplexity) of $y$ under the model $\pi_{\theta}$ as:
$\overline{LP}(y):=\frac{1}{\operatorname{len}(y)}\sum_{t\in y}\bigl{[}-\log P_{\theta}(t\mid\text{context})\bigr{]}.$ (27)
This quantity $\overline{LP}(y)$ measures the average uncertainty or difficulty that the model $\pi_{\theta}$ has in predicting the tokens of $y$. Lower $\overline{LP}(y)$ indicates higher model confidence on a per-token basis.
G.1 Model Uncertainty Reduces with Sequence Length
Conjecture G.1 (Uncertainty Reduction with Sequence Length Assertion (URSLA)).
Consider a set of responses $y$ generated by the model $\pi_{\theta}$. Let $\overline{LP}(y)$ be the average per-token negative log-probability of $y$. The URSLA Conjecture states that for a non-empty subset of responses $\mathcal{Y}$, as the length of a response $y\in\mathcal{Y}$ increases, the expected value of $\overline{LP}(y)$ decreases. Formally, there exists a set $\mathcal{Y}$ such that for all $y\in\mathcal{Y}$:
$\frac{\partial\overline{LP}(y)}{\partial\operatorname{len}(y)}<0,\quad\text{for all }y\in\mathcal{Y}.$ (28)
In other words, longer sequences from $\mathcal{Y}$ are generally easier to predict on a per-token level, yielding lower average negative log-probability than shorter sequences.
This conjecture reflects the empirical observation that models, once committed to a certain semantic or syntactic trajectory in generating text, tend to predict subsequent tokens with increasing ease. We have experimentally verified this phenomenon for the models under consideration.
We note that this conjecture may fail under certain conditions—such as when the model encounters domain shifts, incoherent sequences, or adversarial prompts—where longer sequences do not become inherently easier to predict. Empirical verification and domain-specific analysis are thus required to support or refute this conjecture in practical settings.
G.2 Loss Reduction Reduces Expected Sequence Length
Let $\mathcal{Y}^{-}$ be a set of “negative” responses, as identified by some alignment criterion, where the training objective encourages decreasing their probabilities. Minimizing the objective often translates into increasing $\overline{LP}(y)$ for $y\in\mathcal{Y}^{-}$. Under the URSLA Conjecture G.1, shorter sequences have higher $\overline{LP}(y)$, implying that reducing response lengths for these negative instances is a straightforward way to achieve the desired increase in per-token perplexity and decrease in their probabilities.
Lemma G.2 (Length Reduction for Negative Responses).
Consider a loss function $L(\theta)$ that depends on the probabilities of a set of negative responses $\mathcal{Y}^{-}$. Suppose that reducing the probabilities of these negative responses (equivalently, increasing their average negative log-probability $\overline{LP}(y)$) decreases the overall loss $L(\theta)$. Under Conjecture G.1, there exists a positive correlation between decreasing $\operatorname{len}(y)$ for $y\in\mathcal{Y}^{-}$ and reducing $L(\theta)$. Formally, let:
$\frac{\partial L(\theta)}{\partial\operatorname{len}(y)}>0\quad\text{for all }y\in\mathcal{Y}^{-}.$ (29)
Then, minimizing $L(\theta)$ w.r.t. $\theta$ encourages decreasing $\operatorname{len}(y)$ for $y\in\mathcal{Y}^{-}$.
Proof: Assume that for each $y\in\mathcal{Y}^{-}$, the training objective encourages a reduction in $\pi_{\theta}(y|x)$, the probability assigned to $y$. Equivalently, this corresponds to increasing $-\log\pi_{\theta}(y|x)$, or increasing its average negative log-probability $\overline{LP}(y)$. By the chain rule, we have:
$\frac{\partial L(\theta)}{\partial\operatorname{len}(y)}=\frac{\partial L(\theta)}{\partial\overline{LP}(y)}\cdot\frac{\partial\overline{LP}(y)}{\partial\operatorname{len}(y)}.$ (30)
Since $L(\theta)$ decreases when $\overline{LP}(y)$ increases (to reduce $\overline{\log\pi_{\theta}}(y|x)$), we have:
$\frac{\partial L(\theta)}{\partial\overline{LP}(y)}<0.$ (31)
By the URSLA Conjecture G.1,
$\frac{\partial\overline{LP}(y)}{\partial\operatorname{len}(y)}<0.$ (32)
Combining the two inequalities:
$\frac{\partial L(\theta)}{\partial\operatorname{len}(y)}=\frac{\partial L(\theta)}{\partial\overline{LP}(y)}\cdot\frac{\partial\overline{LP}(y)}{\partial\operatorname{len}(y)}>0,$ (33)
since the product of two negative numbers is positive.
A positive gradient $\frac{\partial L(\theta)}{\partial\operatorname{len}(y)}>0$ implies that reducing $\operatorname{len}(y)$ decreases $L(\theta)$. Hence, during training, the model is incentivized to shorten negative responses $y\in\mathcal{Y}^{-}$. This completes the proof. ∎
G.3 Discussion on the Surprising Relationship between Sequence Length and Average Model (Un)Certainty
Lemma G.2 demonstrates that under the URSLA Conjecture G.1, minimizing a loss function that penalizes high-probability negative responses effectively encourages the model to produce shorter negative responses. While length normalization techniques are often introduced to mitigate trivial solutions (such as producing degenerate short outputs), the above analysis reveals that under certain assumptions about length and certainty, length normalization alone can inadvertently incentivize undesirable length reductions. This necessitates additional considerations, such as length-regularization or more sophisticated constraints, to ensure that the model’s improvements are not merely a byproduct of changing response lengths but reflect genuine quality enhancements.
G.4 Justification for an EOS-Probability Regularizer
Real-world training data often reflect inherent human biases. Humans typically write and record relatively concise content due to time, effort, or domain constraints. However, at inference time, users may request more elaborate, detailed, and longer responses from a model, especially when studying complex topics that require extended reasoning or rich explanations. This discrepancy can create a mismatch between the training distribution and the desired inference-time behavior.
Dataset-Induced Bias for Shorter Sequences.
Human-generated training data often skews toward brevity. Tasks such as writing quick responses, providing minimal clarifications, or answering direct questions do not always require exhaustive elaboration. Economic factors, annotation fatigue, and cognitive load mean human annotators or content creators produce shorter sequences that still suffice to convey basic meaning. As a result, the training corpus is filled with many shorter samples, unintentionally leading the model to treat shorter responses as “normal” or even optimal.
Mismatch Between Training and Inference Requirements.
While training data may be efficient and minimal, users in deployment scenarios often seek more extensive answers. For instance, a user querying a large language model about a technical topic may benefit from a thorough explanation rather than a terse summary. Similarly, users might expect models to produce intermediate reasoning steps, citations, code snippets, or detailed examples. When the model is trained primarily on shorter samples, it fails to internalize the incentive to produce these richer, longer responses at inference time.
Premature EOS as a Shortcut.
In the absence of constraints, a model that has learned from predominantly short examples can exploit early termination as a trivial optimization. By producing an end-of-sequence (EOS) token sooner, it can quickly finalize its response and avoid the complexity of generating lengthy, coherent content. This behavior reduces computational effort and—under certain loss formulations—may even appear optimal, as shorter sequences are well-represented and lower-risk in the model’s learned distribution.
Encouraging Richer Outputs via EOS Probability Control.
To counteract the dataset-induced bias and prevent premature truncation, we introduce a regularizer that influences the probability distribution over the EOS token. By penalizing or rewarding the model for EOS placement, we can realign training incentives with deployment-time requirements. For example, if human evaluators or usage logs indicate a preference for more elaborate answers, we can discourage too-early EOS predictions, nudging the model toward providing fuller, more informative content. Conversely, if the model tends to ramble or produce unnecessarily long responses, the regularizer can be adjusted to encourage timely conclusions.
In essence, the EOS-probability regularizer acts as a corrective mechanism. It helps balance the model’s learned tendency toward short outputs—an artifact of the training data—with the user’s practical need for more extensive, detailed responses at inference time. By bridging the gap between the dataset distribution and real-world user preferences, we ensure that the model’s output quality and length better reflect the desired usage scenarios.
G.5 Incorporating the Regularizer into the Refa Loss
We start from the previously defined length-normalized Refa-dynamic loss (with deviation-based weights and hyperparameters $\alpha$, $\beta$, $\gamma$ as described in Section E). Let:
$\overline{P^{+}}:=\sum_{y\in Y^{+}}e^{\beta(\overline{\log\pi_{\theta}}(y|x)+\alpha\Delta S_{y})},\quad\overline{P^{-}}:=\sum_{y\in Y^{-}}e^{\beta(\overline{\log\pi_{\theta}}(y|x)+\alpha\Delta S_{y})}.$
Then, the Refa-dynamic loss without regularization is:
[…]
To incorporate a preference on the EOS probability, we define a regularization term $\mathcal{R}(\theta)$ that depends on $P_{\theta}(\text{EOS}\mid\text{tokens before EOS in }y)$. If the training data skew toward shorter sequences, we may increase $\lambda$ to discourage prematurely high EOS probabilities, thus encouraging the model to produce more extensive responses. Conversely, if the model generates unnecessarily long and uninformative content, we can adjust the regularizer to incentivize more timely endings.
A simple choice for such a regularizer is:
$\mathcal{R}(\theta)=\sum_{y\in Y^{+}\cup Y^{-}}\lambda P_{\theta}(\text{EOS atposition }|y|),$ (35)
where $\lambda$ is a small constant that scales the influence of this regularizer. A positive $\lambda$ can encourage the model to place appropriate probability mass on the EOS token at the position it appears in the human-written training examples, gradually adjusting the model’s length preferences.
Integrating this regularizer into the loss function, we obtain:
[…]
By tuning $\lambda$ and the form of $\mathcal{R}(\theta)$, practitioners can mitigate the dataset-induced length bias, steering the model toward producing responses of a more desirable length and detail level. This balanced approach ensures that the model’s performance at inference time more closely aligns with user expectations, enhancing both the quality and relevance of its outputs.
Appendix H Experiments Details
H.1 Experimental Setup
Evaluation Benchmarks
We evaluate our models using three widely recognized open-ended instruction-following benchmarks: MT-Bench, AlpacaEval 2, AlpacaEval and Arena-Hard v0.1. These benchmarks are commonly used in the community to assess the conversational versatility of models across a diverse range of queries.
AlpacaEval 2 comprises 805 questions sourced from five datasets, while MT-Bench spans eight categories with a total of 80 questions. The recently introduced Arena-Hard builds upon MT-Bench, featuring 500 well-defined technical problem-solving queries designed to test more advanced capabilities.
We adhere to the evaluation protocols specific to each benchmark when reporting results. For AlpacaEval 2, we provide both the raw win rate (WR) and the length-controlled win rate (LC), with the latter being designed to mitigate the influence of model verbosity. For Arena-Hard, we report the win rate (WR) against a baseline model. For MT-Bench, we present the scores as evaluated by GPT-4-Preview-1106, which serve as the judge model.
Baselines
We compare our approach against several established offline preference optimization methods, summarized in Table . Among these are RRHF (Yuan et al., 2023) and SLiC-HF (Zhao et al., 2023b), which employ ranking loss techniques. RRHF uses a length-normalized log-likelihood function, akin to the reward function utilized by SimPO (Meng et al., 2024), whereas SLiC-HF directly incorporates log-likelihood and includes a supervised fine-tuning (SFT) objective in its training process.
IPO (Azar et al., 2023) presents a theoretically grounded approach that avoids the assumption made by DPO, which treats pairwise preferences as interchangeable with pointwise rewards. CPO (Guo et al., 2024), on the other hand, uses sequence likelihood as a reward signal and trains jointly with an SFT objective.
ORPO (Hong et al., 2024b) introduces a reference-free odds ratio term to directly contrast winning and losing responses using the policy model, also incorporating joint training with the SFT objective. R-DPO (Park et al., 2024) extends DPO by adding a regularization term that mitigates the exploitation of response length.
InfoNCA (Chen et al., 2024), which introduces a K-category cross-entropy loss, reframes generative modeling problems as classification tasks by contrasting multiple data points. It computes soft labels using dataset rewards, applying a softmax operation to map reward values into probability distributions.
Lastly, SimPO (Meng et al., 2024) leverages the average log probability of a sequence as an implicit reward, removing the need for a reference model. It further enhances performance by introducing a target reward margin to the Bradley-Terry objective, significantly improving the algorithm’s effectiveness.